OpenStudy (anonymous):

Solve the system of equations. 2x + 3y = 6 x + y = 4

OpenStudy (anonymous):

\[x+y=4\implies 2x+2y=8\implies y=-2\implies x-2=4\implies x=6\]

OpenStudy (anonymous):

from second equation u get x = 4-y put this is first equation 2(4-y) + 3y = 6 8 - 2y + 3y = 6 y = 6-8 y = -2 Put y = -2 in second equation x + y = 4 x + (-2) = 4 x-2 = 4 x = 4+2 x = 6 solutions (6,-2)